# Multi-Agent Conditional Diffusion Model with Mean Field Communication as Wireless Resource Allocation Planner

### v1

@@ Abstract
In wireless communication systems, efficient and adaptive resource allocation plays a crucial role in enhancing overall Quality of Service (QoS). While centralized Multi-Agent Reinforcement Learning (MARL) frameworks rely on a central coordinator for policy training and resource scheduling, they suffer from scalability issues and privacy risks. In contrast, the Distributed Training with Decentralized Execution (DTDE) paradigm enables distributed learning and decision-making, but it struggles with non-stationarity and limited inter-agent cooperation, which can severely degrade system performance. To overcome these challenges, we propose the Multi-Agent Conditional Diffusion Model Planner (MA‑CDMP) for decentralized communication resource management. Built upon the Model-Based Reinforcement Learning (MBRL) paradigm, MA-CDMP employs Diffusion Models (DMs) to capture environment dynamics and plan future trajectories, while an inverse dynamics model guides action generation, thereby alleviating the sample inefficiency and slow convergence of conventional DTDE methods. Moreover, to approximate large-scale agent interactions, a Mean-Field (MF) mechanism is introduced as an assistance to the classifier in DMs. This design mitigates inter-agent non-stationarity and enhances cooperation with minimal communication overhead in distributed settings. We further theoretically establish an upper bound on the distributional approximation error introduced by the MF-based diffusion generation, guaranteeing convergence stability and reliable modeling of multi-agent stochastic dynamics. Extensive experiments demonstrate that MA‑CDMP consistently outperforms existing MARL baselines in terms of average reward and QoS metrics, showcasing its scalability and practicality for real-world wireless network optimization.
@@ I Introduction
Wirelsess communication networks have been extensively deployed in domains such as smart grids, vehicular communications, and the Internet of Things (IoT) due to their flexibility, low cost, and robustness[1, 2]. Within these networks, efficient resource allocation is vital for optimizing spectrum and energy utilization[3], yet increasing network complexity and the locality of node observations limit system stability and scalability. Traditional rule-based schemes [4, 5] often fail under dynamic or high-load conditions, motivating the use of Reinforcement Learning (RL) for adaptive resource management [6, 7]. However, heterogeneous service demands and the dynamic arrival and departure of services often lead to inefficient resource utilization. In response, Multi-Agent RL (MARL) has emerged to further enhance adaptability by enabling distributed decision-making[8, 9]. Existing MARL frameworks primarily follow either the Centralized Training with Decentralized Execution (CTDE) paradigm or the Distributed Training with Decentralized Execution (DTDE) paradigm[10]. While CTDE improves cooperation and credit assignment through centralized training, it suffers from scalability and privacy issues as the number of agents grows [11]. In contrast, DTDE[12, 13] supports fully distributed learning and local collaboration, offering a more scalable solution for large-scale wireless communication networks.
However, DTDE’s decentralized structure limits global information sharing, which can exacerbate non-stationarity in multi-agent learning[10]. To improve training stability and sample efficiency, we adopt the Model-Based RL (MBRL) paradigm[14], where a world model is learned from historical logs to simulate potential outcomes and generate synthetic trajectories, enabling agents to exploit past experiences while reducing trial-and-error costs[15]. This formulation treats decision-making as a planning problem, facilitating long-term reasoning and informed policy generation[16]. In essence, the world model serves as a generative model that reconstructs future transitions conditioned on current states and actions. Despite efforts like model ensembling [17] and adversarial learning [18], prediction errors in complex and dynamic wireless networks remain a major challenge. Therefore, improving the generative capability of this model is vital for reliable planning. Recent progress in deep generative modeling offers valuable insights. Compared with earlier approaches such as Energy-Based Models (EBMs) [19], Variational Autoencoders (VAEs) [20], and Generative Adversarial Networks (GANs) [21], Diffusion Models (DMs) demonstrate superior sample quality and training stability through an iterative noise-to-data denoising process, representing the forefront of advances in generative learning [22, 23]. These properties make DMs particularly suitable for offline MBRL. Nevertheless, most MARL algorithms incorporating DMs follow the Model-Free RL (MFRL) paradigm [24, 25], where actions are directly generated by diffusion models, often leading to limited sample utilization and unstable convergence. The integration of DMs with multi-agent MBRL thus remains an open and promising direction for further exploration.
@@
On the other hand, another line of difficulty in DTDE is owing to the lack of a centralized coordinator, which makes the design of communication and fusion mechanisms essential yet challenging [26]. A common approach is the consensus-based method[27, 12], which propagates and aggregates information through parameterized functions. While its convergence can be theoretically guaranteed under linear function approximation, it becomes unreliable when extended to neural networks. Another research direction employs Graph Neural Networks (GNNs) to model multi-agent interactions and facilitate cooperation through automated relational learning[28, 29]. Although GNN-based methods can capture complex dependencies, they often suffer from high computational costs and heavy data requirements. Compared to these consensus-based or graph-based methods, Mean Field (MF) [30], which allows local information exchange through approximating all other neighbors’ dynamics, offers both theoretical performance guarantees and computational efficiency.
@@ VI-B1 Performance Comparison
To further evaluate the effectiveness of MA-CDMP, we compare the detailed QoS performance under both ideal and limited RF environments, as illustrated in Fig. 4. In the limited RF scenario, the carrier frequency is increased to introduce higher propagation loss, and the transmission power is reduced to emulate shorter communication ranges with degraded signal quality, following the FSPL model in Eq. (9). As shown in the results, although overall performance slightly declines under the limited RF condition, MA-CDMP consistently achieves higher average throughput, lower average delay, and reduced packet loss rate compared with all baselines. These findings verify that MA-CDMP maintains strong adaptability and robustness across varying channel conditions, underscoring its practicality and reliability in complex and dynamic wireless deployment scenarios.
Figure 4: Comparison of MA-CDMP with other methods in terms of QoS under ideal RF and limited RF scenarios.

### v2

@@ Abstract
In wireless communication systems, efficient and adaptive resource allocation plays a crucial role in enhancing overall Quality of Service (QoS). Compared to the conventional Model-Free Reinforcement Learning (MFRL) scheme, Model-Based RL (MBRL) first learns a generative world model for subsequent planning. The reuse of historical experience in MBRL promises more stable training behavior, yet its deployment in large-scale wireless networks remains challenging due to high-dimensional stochastic dynamics, strong inter-agent cooperation, and communication constraints. To overcome these challenges, we propose the Multi-Agent Conditional Diffusion Model Planner (MA-CDMP) for decentralized communication resource management. Built upon the Distributed Training with Decentralized Execution (DTDE) paradigm, MA-CDMP models each communication node as an autonomous agent and employs Diffusion Models (DMs) to capture and predict environment dynamics. Meanwhile, an inverse dynamics model guides action generation, thereby enhancing sample efficiency and policy scalability. Moreover, to approximate large-scale agent interactions, a Mean-Field (MF) mechanism is introduced as an assistance to the classifier in DMs. This design mitigates inter-agent non-stationarity and enhances cooperation with minimal communication overhead in distributed settings. We further theoretically establish an upper bound on the distributional approximation error introduced by the MF-based diffusion generation, guaranteeing convergence stability and reliable modeling of multi-agent stochastic dynamics. Extensive experiments demonstrate that MA‑CDMP consistently outperforms existing MARL baselines in terms of average reward and QoS metrics, showcasing its scalability and practicality for real-world wireless network optimization.
@@ I Introduction
Wirelsess communication networks have been extensively deployed due to their flexibility, low cost, and robustness[1, 2]. Within these networks, efficient resource allocation is vital for optimizing spectrum and energy utilization[3], yet increasing network complexity and localized observations limit system stability and scalability. Traditional rule-based schemes [4, 5] often fail under dynamic or high-load conditions, motivating the use of Reinforcement Learning (RL) for adaptive resource management [6, 7]. Nonetheless, widely adopted Model-Free RL (MFRL) methods might suffer from unstable training and low sample efficiency, due to their reliance on costly online interactions with the wireless environment. To address this, we adopt the Model-Based RL (MBRL) paradigm[8], where a world model is learned from historical logs to simulate potential outcomes and generate synthetic trajectories, enabling planning-based decision-making with improved long-term reasoning[9, 10].
Despite earlier efforts like model ensembling [11], advanced encoding [12] and adversarial learning [13, 14], building an accurate generative world model in complex and dynamic wireless networks remains a major challenge [15]. Fortunately, compared with Energy-Based Models (EBMs) [16], Variational Autoencoders (VAEs) [12], and Generative Adversarial Networks (GANs) [14], Diffusion Models (DMs) demonstrate superior sample quality and training stability through an iterative noise-to-data denoising process, representing the forefront of advances in generative learning [17, 18]. However, existing works that combine DMs with RL mainly adopt a model-free pipeline [19, 20], where DMs generate actions directly. Although [21] demonstrates the superiority of diffusion-based MBRL for wireless resource allocation, this framework still operates under a centralized paradigm, leading to inefficient resource utilization due to heterogeneous service demands and the dynamic arrival and departure of services in wireless networks. In response, Multi-Agent RL (MARL) has emerged to further enhance adaptability by enabling distributed decision-making[22, 23]. Existing MARL frameworks primarily follow either the Centralized Training with Decentralized Execution (CTDE) paradigm or the Distributed Training with Decentralized Execution (DTDE) paradigm[24]. While CTDE improves cooperation and credit assignment through centralized training, it faces scalability and privacy issues as the number of agents grows [25]. In contrast, DTDE[26, 27] learns fully distributed policies with local collaboration, offering better support for large-scale wireless communication networks.
@@
However, DTDE’s decentralized structure limits global information sharing, which can result in non-stationarity in multi-agent learning[24]. Therefore, the design of communication and fusion mechanisms is essential yet challenging [28]. A common approach is the consensus-based method[29, 26], which propagates and aggregates information through parameterized functions. While its convergence can be theoretically guaranteed under linear function approximation, it becomes unreliable when extended to neural networks. Another research direction employs Graph Neural Networks (GNNs) to model multi-agent interactions and facilitate cooperation through automated relational learning[30, 31]. Although GNN-based methods can capture complex dependencies, they often suffer from high computational costs and heavy data requirements. Compared to these consensus-based or graph-based methods, Mean Field (MF) [32], which allows local information exchange through approximating all other neighbors’ dynamics, offers both theoretical performance guarantees and computational efficiency.
@@ VI-B1 Performance Comparison
To further evaluate the effectiveness of MA-CDMP, we compare the detailed QoS performance under both ideal and limited RF environments, as illustrated in Table III. In the limited RF scenario, the carrier frequency is increased to introduce higher propagation loss, and the transmission power is reduced to emulate shorter communication ranges with degraded signal quality, following the FSPL model in Eq. (9). As shown in the results, although overall performance slightly declines under the limited RF condition, MA-CDMP consistently achieves higher average throughput, lower average delay, and reduced packet loss rate compared with all baselines. These findings verify that MA-CDMP maintains strong adaptability and robustness across varying channel conditions, underscoring its practicality and reliability in complex and dynamic wireless deployment scenarios.
TABLE III: Comparison of MA-CDMP with other methods in terms of QoS under ideal RF and limited RF scenarios.
Algorithms Average Throughput $\uparrow$ Average Delay $\downarrow$ Packet Loss Rate $\downarrow$
Ideal RF Limited RF Ideal RF Limited RF Ideal RF Limited RF
MA-TD3+BC 1638.340 1583.916 0.142 0.197 0.075491 0.105804
MA-$D$ 1653.267 1591.878 0.124 0.180 0.066170 0.096733
MA-Diffuser 1716.524 1662.399 0.065 0.120 0.031228 0.062148
MA-DT 1726.390 1671.691 0.053 0.120 0.025690 0.060918
MA-CQL 1729.809 1673.839 0.051 0.113 0.023786 0.056638
MA-CDMP (Ours) 1732.852 1687.863 0.048 0.102 0.021732 0.049377